OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of a double angle. Then find the exact value of the expression. 2 sin 120° cos 120° is it sqrt -3/2?

OpenStudy (anonymous):

\[\bf 2\sin(120)\cos(120)=\sin(2(120))=\sin(240)=\sin \left( \frac{ 4\pi }{ 3 } \right)=-\frac{ \sqrt3 }{ 2 }\]

OpenStudy (anonymous):

So you're correct =] @SPrezzy

OpenStudy (anonymous):

Thank you!